Astronauts in training are subjected to extreme acceleration forces by the centripetal forces in a giant centrifuge. The radius of the centrifuge is approximately 5 m.




a) Calculate the approximate centripetal force on an astronaut of mass 80 kg if the centrifuge rotates once every 2 s.




b) Approximately how many times greater than the astronauts




weight is this force?

1
Expert's answer
2021-11-18T15:11:24-0500

a)

"a_c=\omega^2 r=\frac{4\pi^2 r}{T^2}=49.3~\frac m{s^2},"

b)

"k=\frac {F_a}P=\frac{a_c}g=5."